### v1

# Time-MMD: A New Multi-Domain Multimodal Dataset for Time Series Analysis

Abstract
Time series data are ubiquitous across a wide range of real-world domains. While real-world time series analysis (TSA) requires human experts to integrate numerical series data with multimodal domain-specific knowledge, most existing TSA models rely solely on numerical data, overlooking the significance of information beyond numerical series. This oversight is due to the untapped potential of textual series data and the absence of a comprehensive, high-quality multimodal dataset. To overcome this obstacle, we introduce Time-MMD, the first multi-domain, multimodal time series dataset covering 9 primary data domains. Time-MMD ensures fine-grained modality alignment, eliminates data contamination, and provides high usability. Additionally, we develop MM-TSFlib, the first multimodal time-series forecasting (TSF) library, seamlessly pipelining multimodal TSF evaluations based on Time-MMD for in-depth analyses. Extensive experiments conducted on Time-MMD through MM-TSFlib demonstrate significant performance enhancements by extending unimodal TSF to multimodality, evidenced by over 15% mean squared error reduction in general, and up to 40% in domains with rich textual data. More importantly, our datasets and library revolutionize broader applications, impacts, research topics to advance TSA. The dataset and library are available at https://github.com/AdityaLab/Time-MMD and https://github.com/AdityaLab/MM-TSFlib.
1 Introduction
Time series (TS) data are ubiquitous across a wide range of domains, including economics, urban computing, and epidemiology (33; 36; 19). Analytical tasks on such datasets hence find broad applications in various real-world scenarios such as energy forecasting, traffic planning, and epidemic policy formulation. Human experts typically complete such Time-Series Analysis (TSA) tasks by integrating multiple modalities of time-series data. For instance, epidemiologists combine numerical data on influenza infections with textual domain knowledge, policies, and reports to predict future epidemiological trends. However, most existing TSA models (38; 26; 54; 28; 56; 20; 55; 30) are unimodal, solely using numerical series.
Recently, with the development of Large Language Models (LLMs), the field of TSA is also undergoing an exciting transformative moment with the integration of natural language (50; 21). Existing LLM-based TSA methods incorporate endogenous text derived from numerical series, such as linguistic descriptions of statistical information, which has demonstrated promising benefits (15; 18; 4; 24). However, the potential of exogenous or auxiliary textual signals—such as information on concurrent events and policies that provide additional context to time series—remains untapped. This observation prompts a crucial question for multimodal TSA: Can multimodal TSA models utilize these exogenous textual signals effectively, thereby enhancing current TSA tasks and enabling new applications?
The primary obstacle in addressing this question lies in the absence of a comprehensive, high-quality multimodal TS dataset, as evidenced by three significant gaps: (1) Narrow data domains. Data characteristics and patterns vary between different domains, such as the periodicity of numerical data and the sparsity of textual data. However, current multimodal TS datasets (12; 48; 9; 31; 4) focus solely on stock prediction tasks in the financial domain, which are unable to represent the diverse data domains. (2) Coarse-grained modality alignment. Existing multimodal TS datasets only ensure that the text and numerical data come from the same domain, such as general stock news and the prices of one specific stock. Clearly, an abundance of irrelevant text diminishes the effectiveness of multimodal TSA. (3) Inherent data contamination. Existing multimodal TS datasets overlook two main reasons of data contamination: (1) Textual data often contains predictions. For example, influenza outlook is a regular section in influenza reports. (2) Outdated test set, particularly the textual data, may have been exposed to LLMs, which are pretrained on vast corpuses. For example, the knowledge cutoff for Llama3-70B is December 2023, which is later than the cutoff dates for most existing multimodal TS datasets. These reasons lead to biased evaluations of general or LLM-based TSA models.
To address the identified gaps, this work aims to introduce a comprehensive, high-quality multimodal TS dataset that spans diverse domains and can be validated through its effectiveness and benefits for TSA. The main contributions of our work are:
Pioneering Multi-Domain Multimodal Time-Series Dataset: Time-MMD. We introduce Time-MMD, the first multi-domain multimodal time-series dataset that addresses the aforementioned gaps: (1) encompasses 9 primary data domains. (2) ensures fine-grained modality alignment through meticulously selected data sources and rigorous filtering steps. (3) disentangles facts and predictions from text; ensures all cutoff dates are up to May 2024. To the best of our knowledge, Time-MMD stands as the inaugural high-quality and comprehensive multimodal time-series dataset. We envision Time-MMD offering exciting opportunities to significantly advance time series analysis through multimodal extensions.
Pilot Multimodal Time-Series Forecasting Library: MM-TSFlib. We develop the first multimodal time-series forecasting (TSF) library, MM-TSFlib, piloting multimodal TSA research based on Time-MMD. Our library MM-TSFlib features an end-to-end pipeline with a seamless interface that allows the integration of any open-source language models with arbitrary TSF models, thereby enabling multimodal TSF tasks. MM-TSFlib facilitates easy exploration of Time-MMD and supports future advancements in multimodal TSA.
Extensive Evaluations with Significant Improvement. We conducted experiments of multimodal TSF on Time-MMD using MM-TSFlib. The multimodal versions outperformed corresponding unimodal versions in all TSF backbones, reducing the mean squared error by an average of over 15% and up to 40% in some domains with rich textual data. This significant and consistent improvement demonstrates the high quality of Time-MMD, the effectiveness of MM-TSFlib, and the superiority of multimodal extensions for TSF.
We include additional related works in Appendix A and limitations in Appendix B.
2 Multi-Domain Multimodal Time-Series Dataset: Time-MMD
We first introduce the key challenges in constructing Time-MMD, followed by the construction pipeline. We then detail each component of the pipeline with corresponding data quality verification. Finally, we discuss considerations for fairness and data release.
Challenges. Creating a high-quality, multi-domain numerical-text series dataset presents significant challenges, encompassing the effective gathering, filtering, and alignment of useful textual data. First, textual sources are sparse. Unlike numerical data, typically provided by a "packaged" source, textual data are collected from a variety of dispersed sources, such as reports and news articles, necessitating extensive individual collection efforts. Second, textual information is noisy. Raw textual data often contains large portions of irrelevant information and potential data contamination, such as expert predictions in reports, requiring rigorous filtering processes to ensure data quality. Third, textual data requires precise alignment. It is essential to achieve temporal alignment between textual and numerical data by synchronizing reported times with numerical time steps (e.g., the time step where text is posted) and ensuring that the effective duration of textual information matches the relevant time frames at various granularities (e.g., a seasonal report should correspond to 12 time steps in a weekly time series). Additionally, the dataset faces challenges regarding ease of use, maintenance, and regular updates to remain relevant and useful for ongoing research and applications.
Pipeline Overview. We propose a comprehensive pipeline for constructing a text-numeric series dataset utilizing modern LLMs. As illustrated in Figure 1, the construction process is divided into three key steps: (1) Numerical Series Data Construction. We gather numerical data from reputable sources to ensure reliability and accuracy. (2) Textual Series Data Construction. Textual data is collected for fine-grained matching with the numerical data. The quality of this matching is ensured through human selection of data sources and raw text filtering by LLMs. Additionally, LLMs are employed to disentangle facts and predictions and generate summaries. (3) Numerical-Textual Alignment. We use binary timestamps to mark the start and end dates as a universal temporal alignment method between numerical and textual series, supporting the requirements of various downstream TSA tasks.
2.1 Numerical Series Data Construction
Data Source Selection. We select data sources that are (1) reliable, containing verified knowledge; (2) actively released, allowing for updates with new data; and (3) multi-domain, covering various TSA patterns. Based on these principles, we choose nine data sources from different domains. Most sources are from government agencies, with the lowest update frequency being six months.
Target Variable Selection. For each domain, we select target variables with significant real-world implications, indicating easier text matching, as shown in Table 1. These variables span three distinct frequencies: daily, weekly, and monthly.
Collection & Preprocessing. We collect raw data for all available times, either from batch-released files or through individual scraping. We preprocess the data by discarding early years with a high proportion of missing values. We maintain the original frequency for most domains, adjusting it for security and climate domains due to irregular releases and difficult text matching, respectively. Figure 2 illustrates the diverse patterns present in each domain, such as periodicity and trends.
Data Quality & Property. As shown in Table 1 and Figure 2, the constructed numerical data provides comprehensive temporal coverage, ranging from the earliest in 1950 to the present, and exhibits distinct patterns, such as periodicity and trends.
| Domain | Target | Frequency | Timestamps | Timespan |
| Agriculture | Retail Broiler Composite | Monthly | 1983 - Present | |
| Climate | Drought Level | Monthly | 1983 - Present | |
| Economy | International Trade Balance | Monthly | 1989 - Present | |
| Energy | Gasoline Prices | Weekly | 1996 - Present | |
| Environment | Air Quality Index | Daily | 1982 - 2023 | |
| Health | Influenza Patients Proportion | Weekly | 1997 - Present | |
| Security | Disaster and Emergency Grants | Monthly | 1999 - Present | |
| Social Good | Unemployment Rate | Monthly | 1950 - Present | |
| Traffic | Travel Volulmn | Monthly | 1980 - Present |
2.2 Textual Series Data Construction
Data Source Selection: Selected Reports and Web Search Results. The choice of data sources should take into account both extensive coverage and initial strong relevance to the numerical data. Consequently, we combine two appropriate data source types as follows: (1) Selected Reports: For each target variable, we manually select 1-2 highly relevant report series with guaranteed updates. For instance, the weekly influenza report111https://www.cdc.gov/flu/weekly/weeklyarchives2023-2024/week04.htm published by the Centers for Disease Control and Prevention of the United States is chosen as one of the report sources for the weekly influenza patients proportion of the United States. (2) Web Search Results: For each target variable, we design 2-3 highly relevant keywords used for web searching.
These two data sources complement each other: report data ensures higher relevance but cannot guarantee all-time coverage, while search results cover all times but are highly redundant; search results aggregate multiple data sources, while report data, usually in PDF or TXT format, cannot be extracted by searching.
Data Collection: Searching and Crawling. For keyword web searching, we use the official Google API222https://developers.google.com/custom-search/v1/overview as the entry point. For each keyword, we collect the timestamp, source, title, and content from the top 10 results located each week from 1980 to present. For report data, we parse all available reports from each data source and preserve only plain-text paragraphs.
Data Preprocessing: Filtering, Disentangling, and Summarizing. To curate the collected raw text data, we introduce three key preprocessing steps: (1) Filtering to improve relevance; (2) Disentangling facts with predictions to mitigate data contamination; (3) Summarizing for better usability. Given the impracticality of performing these steps manually, we leverage the state-of-the-art LLM, Llama3-70B, to accomplish these tasks.
The prompt used for LLMs is detailed in Appendix D. We incorporate three specific strategies to alleviate the hallucination issue in LLMs and enhance preprocessing quality: (1) A concise introduction of the text. (2) Mandating the LLM to reference the data source, aiding constraint and verification. (3) Permitting the LLM to indicate ‘not available’ when relevance is uncertain, to avoid fabrication. Appendix E provides a showcase of the text before and after processing.
Data Quality & Property. Overall, Figure 3 visualizes the extracted fact count per month over time by domain. Note that the Agriculture report data is of high volume around 2020 and therefore produces a peak. We make the following observations: (a) The search data count exhibits a gradual increasing trend, benefiting from the development of the Internet; the report data count has stabilized in recent years, indicating that release schedule has become stable. (b) The sparsity of textual data varies across different domains, with high-profile fields often accompanied by richer textual data. These validate the extensibility and updatability of Time-MMD and highlight the importance of its coverage across 9 diverse domains.
We further validate the effectiveness of key steps in textual data construction:
(a) Data sources selection. We use relevance and coverage ratio to describe the percentage of relevant texts and the proportion of numerical series data being covered by at least one fact, respectively. As demonstrated by Table 2, report data exhibits higher relevance but lower coverage; search data display the opposite pattern. Thus, our combined usage serves as a comprehensive solution.
(b) Data preprocessing. Figure 4 provides word cloud visualizations of constructed text data in the health domain, respectively for extracted facts, extracted predictions, and discarded text. Recall that the target variable here is the influenza patients proportion. Highly relevant words such as "pandemic", "vaccine", and "flu" appear more frequently in the extracted facts; research paper-related words such as "edu", "mdpi", and "university" are more common in the discarded text. Besides, the prediction text primarily contains words describing future, such as "will" and "next". These validate the effectiveness of LLM filtering and disentangling. Furthermore, Table 2 presents a comparison of the token count before and after preprocessing. The substantial decrease validates that LLM summarizing improves usability. Appendix F provides the manual verification results on a subset of the data to further validate the effectiveness of preprocessing using LLMs.
| Source | Raw Tokens | Preprocessed Extracted Facts | Preprocessed Extracted Prediction | ||||
| Relevance(%) | Coverage (%) | Tokens | Relevance(%) | Coverage (%) | Tokens | ||
| Report | k | ||||||
| Search | |||||||
2.3 Binary Time Stamps for Diverse TSA Tasks
To enable the Time-MMD for versatile and flexible use, we maintain binary timestamps for all numerical and text data, storing the manually verified start dates and end dates. Such binary stamps can be easily referred to while serving different tasks. For report text data, we manually verify the timestamps based on the release notes or report contents. For search data, we integrate adjacent search results within each week and mark timestamps correspondingly.
2.4 Considerations for Fairness and Data Release
To consider fairness, we gather data from both the United States and African regions in the Health domain. As depicted in Figure 2, the numerical data of African region exhibits weaker periodicity. Figure 3 shows that the African region has considerably fewer reports compared to the United States. We urge researchers to consider underrepresented groups when conducting multimodal TSA tasks.
To support various existing and potential novel TSA tasks, we include the following metadata when releasing Time-MMD: (1) Numerical Data: start & end time, target variable, other variables; (2) Text Data: start & end time, fact text (content & data source), prediction text (content & data source).
3 Multimodal Time-Series Forecasting Library: MM-TSFlib
In this section, we aim to illustrate the potential benefits of our Time-MMD for multimodal TSA by focusing on time-series forecasting (TSF), a fundamental TSA task. TSF involves predicting future events or trends based on historical time-series data. While most existing TSF methods primarily depend on numerical series, we aim to extend these unimodal TSF methods to multimodality. To achieve this, we contribute both formulating the multimodal TSF problem as well as introducing MM-TSFlib, the first comprehensive multimodal TSF library.
3.1 Problem Formulation
Conventional unimodal TSF models take a numerical series as input and output future values of some or all of its features. Let the input variable of the numerical series be denoted as , where is the length of the lookback window decided by domain experts and is the feature dimension at each time step. The output variable of the forecasts generated of horizon window length is denoted as , where is the dimension of targets at each time step. For the sample at time step , denoted as , and . Thus, the unimodal TSF model parameterized by is denoted as .
For multimodal TSF, the input variable of the textual series is also considered, which can be denoted as , where is the lookback window length of the text series, independent of , and is the feature dimension of the text. Although the text variable may have inconsistent feature dimensions, we slightly abuse the notation here for brevity. Thus, the multimodal TSF model parameterized by is denoted as .
3.2 Pioneering Solution for Multimodal TSF
Multimodal Integration Framework. We propose a pioneering multimodal integration framework to extend existing unimodal TSF models to their multimodal versions. As illustrated in Figure 6, our framework features an end-to-end pipeline that integrates open-source language models with various TSF models. Numerical and textual series are independently modeled using unimodal TSF models and LLMs models with projection layers. These outputs are then combined using a learnable linear weighting mechanism to produce the final prediction. To reduce computational costs, we freeze the LLM parameters and train only the additional projection layers. We employ pooling layers to address the inconsistent dimensions of textual variables. This framework features an end-to-end training manner, with marginal trainable parameters overheads.
Multimodal TSF Library. Building upon the multimodal dataset Time-MMD and integration framework, we present the first multimodal TSF library, named MM-TSFlib. MM-TSFlib supports multimodal extensions of over 20 unimodal TSF algorithms through 7 open-source (large) language models, including BERT(11), GPT-2(32) (Small, Medium, Large, Extra-Large), Llama-2-7B(37), and Llama-3-8B333https://llama.meta.com/llama3. We detail the implementations and language models in Appendix G.
MM-TSFlib is designed for ease of use with Time-MMD in multimodal TSA. Additionally, MM-TSFlib serves as a pilot toolkit for evaluating the multimodal extensibility of existing TSF models.
4 Experiments for Multimodal TSF
Based on the constructed MM-TSFlib, we further conduct comprehensive experiments to demonstrate the superiority of multimodal TSF and the high quality of Time-MMD.
4.1 Experimental Setup
We adhere to the general setups following existing TSF literatures(45; 44; 30). Regarding the horizon window length, we consider a wider range from short to long-term TSF tasks, with four different lengths for each dataset according to frequency. We conduct TSF tasks on all 9 domains of Time-MMD. We employ the widely-adopted mean squared error (MSE) as the evaluation metric.
We comprehensively consider 12 advanced unimodal TSF methods across 4 types including: (1) Transformer-based. We deployed two sets of experiments upon TSF models: Transformer (38), Reformer (20), Informer (55), Autoformer (45), Crossformer (54), Non-stationary Transformer (28),FEDformer (57), iTransformer (26). (2) MLP-based: DLinear (51). (3) Agnostic: FiLM (56). (4) LLM-based: Time-LLM(18). Unless otherwise specified, we use GPT-2-Small as the LLM backbone in MM-TSFlib. More details about setup are provided in Appendix H.
unclear correlation between multimodal TSF performance and LLM natural language capabilities
4.2 Experimental Results
Our experiments aim to investigate the following five aspects.
Effectiveness of multimodal TSF. Figure 6 shows average MSE results for corresponding unimodal and multimodal versions of each TSF backbone. The multimodal versions consistently outperform corresponding unimodal versions, reducing the mean squared error by over 15% in average and up to 40% in domains with rich textual data. Such significant improvements fully validate the superiority of multimodal TSF and our multimodal framework in Section 3.
Additionally, we observe that different TSF backbones benefit from multimodal extension to varying degrees. For example, the originally inferior Informer exhibits strong multimodal performance, which we attribute to its intrinsic design for modeling long-range dependencies that may benefit more from textual cues. We hope these results inspire more advanced multimodal TSF models.
Quality of Time-MMD dataset. Figure 6 shows that SOTA unimodal TSF models, such as iTransformer and PatchTST, maintain leading unimodal performance, validating the quality of Time-MMD’s numerical data. Moreover, multimodal extension consistently and significantly improves performance by incorporating textual data, confirming the quality of Time-MMD’s textual data
Influence of data domains. Figure 7a shows the relationship between the relevant fact count and the reduced MSE via multimodal extension for each domain. The scatter plot generally illustrates a positive linear correlation, aligning with the innovation of integrating textual information. Besides, domain characteristics also influence multimodal performance, even with a similar fact count. For example, the security domain, focusing on disasters and emergency grants, exhibits higher unpredictability in the future thus benefits less from the historical textual information. This observation highlights the importance of Time-MMD’s coverage of 9 domains.
Influence of the LLM backbone. We varied the LLM backbone in MM-TSFlib and evaluated corresponding multimodal performance on health domain. As shown in Figure 7b, the choice of LLM backbone does not exhibit a significant correlation with multimodal TSF performance. For the GPT2 series, the scaling law is unclear for multimodal TSF, indicating no clear positive correlation between the parameter scale and TSF performance. Across different LLMs, multimodal TSF performance is relatively similar, even between the advanced Llama-3-8B and the earlier BERT. There might be two possible reasons: (1) Our proposed multimodal framework, although effective, still does not fully utilize the power of LLMs, particularly by only fine-tuning through projection layers. (2) Existing LLMs, pre-trained for natural language tasks, may not be directly suitable for multimodal TSF. Either reason suggests that there is significant room for improvement in multimodal TSF.
Influence of the horizon window size. Figure 7c shows the relationship between horizon window size and the average MSE reduction for each domain. Overall, the MSE reduction is stable and promising across different horizon window size, from short term to long term. This demonstrates that the effectiveness of multimodal TSF is robust to different forecast horizon requirements.
5 Broader Impacts
Beyond its efficacy in enhancing time-series forecasting accuracy through multimodality (Section 4), Time-MMD holds significant potential in advancing time series analysis across a wide spectrum. In the following section, we discuss how Time-MMD can transform conventional approaches, facilitate novel methodologies, and broadly impact the time series analysis domain.
Multimodal Time-Series Imputation. Missing values in time series data, caused by sensor failures, system instability, or privacy concerns, pose a significant challenge in analysis. Conventional time-series imputation (TSI) methods (6; 14; 1; 2) often overlook valuable information captured in textual formats alongside the numerical data. For instance, incident reports, weather conditions, and special events can provide crucial context for imputing missing data points in traffic time series, but current methods fail to effectively incorporate this information. Time-MMD enables the integration of textual contextual information with numerical time series data, opening new avenues for multimodal time series analysis and enhancing imputation accuracy.
Multimodal Time-Series Anomaly Detection. Detecting anomalies in time series data is crucial for identifying unusual patterns that may indicate faults, fraud, or other significant events (8; 3; 52). However, conventional anomaly detection methods (29; 5; 47; 34) are limited to expected pattern deviations in numerical data, overlooking valuable information in textual formats. For instance, news articles, social media posts, and market reports can provide critical context that influences financial market behavior and helps identify anomalies not evident from numerical data alone.
Multimodal Foundation Time-Series Models The introduction of Time-MMD, a comprehensive text-numeric TS dataset, is expected to significantly advance multimodal-based TS methods, including the the development of multimodal foundation TS models. Time-MMD will facilitate further exploration with more complex and informative prompts, enhancing the performance and capabilities of fine-tuning methods. Additionally, it will spur research and development of multimodal models specifically tailored for TSA, an area with limited exploration compared to other domains like vision and video generation(35; 41; 46; 13).
6 Ethics Statement
While collecting data from government and news websites, we rigorously adhered to ethical standards to ensure compliance with website policies and avoid potential conflicts of interest. Mindful of copyright and regional policies, we restricted our collection to content freely available without premium access or subscription requirements. In collecting data from web searches, we used Google’s official API to ensure that the data strictly complied with ethical standards.
7 Conclusion
In this work, we propose Time-MMD, the first multi-domain multimodal time series dataset, and develop MM-TSFlib, the first multimodal time series forecasting library, which facilitates a pilot study for multimodal time series analysis on Time-MMD. We conduct extensive experiments to demonstrate the high quality of Time-MMD, the effectiveness of MM-TSFlib, and the superiority of integrating textual information for time series analysis. We envision that this work catalyzes the transformation of time series analysis from unimodal to multimodal by integrating natural language.
Acknowledgements
This paper was supported in part by the NSF (Expeditions CCF-1918770, CAREER IIS-2028586, Medium IIS-1955883, Medium IIS-2106961, PIPP CCF-2200269, IIS-2008334, CAREER IIS-2144338), CDC MInD program, Meta faculty gifts, and funds/computing resources from Georgia Tech.
Appendix
Appendix A Additional Related Dataset Work
Additional multi-modal datasets[17] for news impact prediction and electrocardiogram classification[21] are proposed recently. However, these datasets still do not address the aforementioned gaps. Especially constructing multimodal datasets for other domains, such as agriculture and security, is more challenging but holds substantial real-world impact.
Appendix B Limitations
Our work provides a comprehensive, high-quality multimodal time series dataset, but it still has limitations in terms of dataset diversity, as all the text data comes from English. We plan to extend Time-MMD to multilingual versions to better address diversity and leverage data from multiple languages. Furthermore, our multimodal time series forecasting library is built upon a simple integration framework that only uses a projection layer for fine-tuning LLMs. How to fine-tune LLMs more efficiently and effectively for time series analysis remains an interesting topic. Despite these limitations, we hope that the datasets and library we have constructed will facilitate broader research and applications in multimodal time series analysis.
Appendix C Statistics of Textual Data
We show statistics of collected report data and search data in Table 3 and Table 4. As the security reports themselves contain manually-written summaries, we didn’t perform a LLM preprocessing on them.
| Domain | Raw Tokens | Fact | Prediction | ||||
| Relevance(%) | Coverage(%) | Tokens | Relevance(%) | Coverage(%) | Tokens | ||
| Agriculture | |||||||
| Climate | |||||||
| Economy | |||||||
| Energy | |||||||
| Environment | |||||||
| Health (US) | |||||||
| Health (AFR) | |||||||
| Security | - | - | - | - | - | - | |
| Social Good | |||||||
| Traffic | |||||||
| Domain | Raw Tokens | Fact | Prediction | ||||
| Relevance(%) | Coverage(%) | Tokens | Relevance(%) | Coverage(%) | Tokens | ||
| Agriculture | |||||||
| Climate | |||||||
| Economy | |||||||
| Energy | |||||||
| Environment | |||||||
| Health(US) | |||||||
| Health(AFR) | |||||||
| Security | |||||||
| SocialGood | |||||||
| Traffic | |||||||
Appendix D Prompt designed for LLM preprocessing
We show the prompt we use while doing LLM preprocessing in Figure 8. Corresponding keywords come from manually written domain-specific instructions.
Appendix E Showcase of Raw and Preprocessed Text
We show the pre and post-processing text content of a report in Climate domain in Figure 9. We see LLM accurately summarized the factual content from original report.
Appendix F Manual Verification of LLM Preprocessing
To validate the quality of LLM preprocessing, we manually inspected 100 text samples. Our observations are as follows:
Among the 127 extracted facts, 8 were fabricated by the LLM.
In the 8 fabricated instances, 5 were labeled with the data source as LLM, indicating that they can be filtered out. his results in a true hallucination rate of 3/127.
Among the 52 text samples discarded by the LLM, 4 were manually identified as containing relevant information, yielding an error discard rate of 4/52.
Appendix G Details of MM-TSFlib
In terms of implementation, MM-TSFlib chooses to extend the widely used Time-Series Library (TSlib)444https://github.com/thuml/Time-Series-Library, thus ensuring ease of use. For LLM invocation, MM-TSFlib utilizes the popular and active Hugging Face555https://huggingface.co/models. For the projection layer, we use a multilayer perceptron (MLP) to keep it simple. To use Time-MMD dataset for TSF task, MM-TSFlib constrains the latest end date of the input text sequence to be earlier than the latest end date of the input sequence, in order to avoid information leakage. Overall, MM-TSFlib supports multimodal extensions of over 20 unimodal TSF algorithms via 7 open-source LLM models.
G.1 List of Supported TSF
So far, our models supports following TSF models: TimeMixer [40], TSMixer [7], iTransformer [26], PatchTST [30], TimesNet [44], DLinear [51], LightTS [53], ETSformer [43], Non-stationary Transformer [28], FEDformer [57], Pyraformer [25], Autoformer [45], Informer [55], Reformer [20], Transformer [38], Mamba [16], SegRNN [23], Koopa [27], FreTS [49], TiDE [10], FiLM [56], MICN [39], Crossformer [54], TFT [22].
G.2 List of Supported LLM
For LLM, MM-TSFlib supports BERT[11], GPT-2[32] (Small, Medium, Large, Extra-Large), Llama-2-7B[37], and Llama-3-8B666https://llama.meta.com/llama3.
Appendix H More Details of Experimental Setup
H.1 Time-Series Forecasting Backbones
We deployed two sets of experiments upon TSF models: (1) Transformer-based, including Transformer [38], Reformer [20], Informer [55], Autoformer [45], Crossformer [54], Non-stationary Transformer [28],FEDformer [57], iTransformer [26]. (2) MLP-based: DLinear [51]. (3) Agnostic: FiLM [56]. (4) LLM-based: Time-LLM[18].
Transformer, a classic sequence-to-sequence model basing on multi-head attention mechanism.
Reformer, an computational-efficient Transformer with advancements in attention hashing and reversible residual layers
Informer, an advanced Transformer designed to tackle long-term forecasting problem with sparse attention layers and self-attention distilling.
Autoformer, a Transformer-based model that keeps encoder-decoder structure but alters attention computations by auto-correlation mechanism in order to benefit long-term forecasting.
Crossformer, a multi-variate Transformer-based model that explicitly explores and utilizes cross-dimension dependencies.
Non-stationary Transformer, a Transformer that is designed to capture non-stationarity patterns instead of temporal correlation.
FEDformer, a Transformer that explicitly use Fourier decomposition results to enhance long-term forecasting ability.
iTransformer, a inverted Transformer that tokenizes multivariate time-series upon each timestamps/
DLibear, a linear model that performs forecasting by a direct regression upon historical time series with a one-layer linear model.
FiLM, a model-agnostic method that introduces Legendre and Fourier projections to denoise series and approximate historical information.
Time-LLM, a framework that integrates LLM for time-series forecasting by reprogramming input series and then aligning with text prototypes.
H.2 LLM Backbones
We use GPT-2 Small [32] for the majority of experiments, while other GPT-2 models and BERT, Llama-2, and Llama-3 are used in the ablation study.
BERT a bidirectional transformer pre-trained on large text corpora with tasks like masked language modeling and next sentence prediction
GPT-2 an advanced language model that generates coherent and contextually relevant text by predicting subsequent words in a sentence, pre-trained on diverse internet text and capable of performing a variety of language tasks without task-specific fine-tuning
Llama-2 an accessible, open-source LLM designed to generate coherent and contextually relevant text by leveraging advanced transformer-based architecture
Llama-3 the latest iteration of the Llama model, offering enhanced text generation and comprehension abilities, further advancing the performance and versatility of its predecessors
H.3 Evaluation Metrics
We use mean squared error (MSE) as the evaluation metric for all experiments. The MSE is defined as the average of the squares of the errors. The error here is the difference between the forecasted value and the actual value . That is:
H.4 Details of Implementation
All experiments are conducted using Pytorch on a 80GB NVIDIA A100 GPU. We follow the commonly adopted setup for defining the forecasting horizon window length, as outlined in prior works[45, 44, 30]. Specifically, for daily reported datasets, the forecasting horizon windows are chosen from the set [48, 96, 192, 336], with a fixed lookback window size of 96 and a consistent label window size of 48 for the decoder. Similarly, for the weekly reported dataset, we employ forecasting horizon windows from [12, 24, 36, 48], with a fixed lookback window size of 36 and a constant label window size of 18 for the decoder. Besides, for the monthly reported dataset, we employ forecasting horizon windows from [6,8,10,12], with a fixed lookback window size of 8 and a constant label window size of 4 for the decoder. When extending unimodal TSF methods to their multimodal versions, we keep the TSF model architecture and parameters unchanged to ensure a fair comparison.